Determine if the following are probability distributions (if no, state why).





a. x 3 6 9 12 15



p(x) 4/10 2/10 1/10 1/10 1/10




b. x 1 2 3 4 5



p(x) 3/15 1/15 1/15 4/15 6/15




c. x 20 30 40 50



p(x) 1.1 0.2 0.9 0.3

1
Expert's answer
2022-03-11T05:55:15-0500

Solution

a.      ∑p(x)= 0.4+0.2+0.1+0.1+0.1

= 0.9

 it’s not a probability distribution

b.      ∑p(x)= 3/15 + 1/15 + 1/15 + 4/15 + 6/15

= 1

It is a distribution function

c.      ∑p(x)= 1.1+0.2+0.9+0.3

= 2.5

It is not a probability distribution